OpenStudy (anonymous):

What is the cube root of 27x18? 3x6 3x9 9x6 9x9

OpenStudy (anonymous):

@AlexandervonHumboldt2 @Nnesha @FLAMER123

OpenStudy (anonymous):

first you must put it \[\sqrt{27.18}\] =\[\sqrt{9*3*9*2}\]=\[9\sqrt{5}\]

OpenStudy (anonymous):

Wait I mean \[9\]

OpenStudy (anonymous):

\[9\sqrt{6}\]

OpenStudy (anonymous):

c is answer

OpenStudy (alexandervonhumboldt2):

C

OpenStudy (michele_laino):

@FLAMER123 sorry, is a cube root not a square root!

OpenStudy (anonymous):

\[\left( 27\times18 \right)^{\frac{ 1 }{ 3 }}\] \[=\left( 3^3 \right)^{\frac{ 1 }{ 3 }}\left( 18 \right)^{\frac{ 1 }{ 3 }}=3\left( 18 \right)^{\frac{ 1 }{ 3 }}\]